OpenStudy (anonymous):

Help please??? What is the equation of a circle with a center at (0, 9) and a radius of 9?

OpenStudy (anonymous):

(x-0)² + (y-9)² = 9²

OpenStudy (anonymous):

then you need to simplify

jimthompson5910 (jim_thompson5910):

Similar Example: Let's say we ask "what is the equation of the circle with radius 4 and center (-2,3)?" Center: \(\large ({\color{red}{h}},{\color{blue}{k}}) = ({\color{red}{-2}},{\color{blue}{3}})\) Radius: \(\large r = {\color{green}{4}}\) Plug all this into \[\Large (x-{\color{red}{h}})^2+(y-{\color{blue}{k}})^2 = {\color{green}{r}}^2\] to get \[\Large (x-{\color{red}{h}})^2+(y-{\color{blue}{k}})^2 = {\color{green}{r}}^2\] \[\Large (x-({\color{red}{-2}}))^2+(y-{\color{blue}{3}})^2 = {\color{green}{4}}^2\] \[\Large (x+2)^2+(y-3)^2 = 16\]

OpenStudy (anonymous):

okay so the answer is x^2+(y-9)^2=9^2?

jimthompson5910 (jim_thompson5910):

you can simplify the 9^2 part

OpenStudy (anonymous):

okay so x^2+(y-9)^2=81?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Good job do you understand though?